# E(n) Equivariant Message Passing Cellular Networks

### v1

@@ 5.1 N-Body System
As introduced by Kipf et al. (2018), the N-body system experiment involves tracking the trajectories of five charged particles in three-dimensional space over time. The objective is to predict the positions of all particles after 1,000 time steps, starting from their initial positions and velocities. To ensure a consistent comparison, we adopted the experimental setup as described by Satorras et al. (2021). While the system operates according to physics, with particles being attracted or repelled based on their charges, the task remains equivariant to rotation and translations, maintaining these symmetries throughout the entire trajectory.
@@
Task $\alpha$ $\Delta\varepsilon$ $\varepsilon_{\text{HOMO}}$ $\varepsilon_{\text{LUMO}}$ $\mu$ $C_{v}$ $G$ $H$ $R^{2}$ $U$ $U_{0}$ ZPVE Units (bohr3) (meV) (meV) (meV) (D) (cal/mol K) (meV) (meV) (bohr3) (meV) (meV) (meV) NMP .092 69 43 38 .030 .040 19 17 .180 20 20 1.50 SchNet * .235 63 41 34 .033 .033 14 14 .073 19 14 1.70 Cormorant .085 61 34 38 .038 .026 20 21 .961 21 22 2.02 TFN .223 58 40 38 .064 .101 - - - - - - SE(3)-Tr. .142 53 35 33 .051 .054 - - - - - - DimeNet++ * .043 32 24 19 .029 .023 7 6 .331 6 6 1.21 SphereNet * .046 32 23 18 .026 .021 8 6 .292 7 6 1.21 PaiNN * .045 45 27 20 .012 .024 7 6 .066 5 5 1.12 SEGNN .060 42 24 21 .023 .031 15 16 .660 12 15 1.62 MPSN .266 153 89 77 .101 .122 31 32 .887 33 33 3.02 EMPSN .066 37 25 20 .023 .024 6 9 .101 7 10 1.37 EGNN .071 48 29 25 .029 .031 12 12 .106 12 12 1.55 Decoupled EMPCN .063 40 27 22 .026 .027 10 10 .104 10 11 1.51 Gain 11% 17% 7% 12% 10% 13% 17% 17% 2% 17% 9% 3%
@@ 5.2 QM9
As indicated in Table 2, incorporating this additional ring information yielded an average improvement of approximately 10% over the baseline EGNN model. While one could potentially generate additional rings via the RadiusGraph or Vietoris-Rips complex, our experiment demonstrates that even with the initial graph’s topology, we can achieve significant performance gain without extra computational complexity.
@@ 5.4 Ablation Study
These findings highlight the robustness and efficiency of decoupled EMPCNs, illustrating that they are a more reliable choice for scenarios where data and computational resources are limited as is typical in real-life settings.
@@ 6 Conclusion
In summary, EMPCNs provide a scalable and expressive framework for higher-order message passing in geometric and topological graphs. Although this method is proposed for cellular complexes, the same techniques could be adapted to other higher-order path message passing methods, potentially enhancing the expressiveness by incorporating various structures beneficial for specific tasks. We believe that exploring the intersection between geometric and topological MPNNs is a crucial aspect of developing efficient and strongly generalizing models, especially in low-data settings.
@@
The main limitation of EMPCNs is the computational cost associated with the calculation of the introduced geometric invariants. Although it is possible to retain only those invariants relevant to a specific downstream task, an interesting direction for future research would be to pre-aggregate cell representations and compute the geometric invariants solely for key structures essential to the task. For the decoupled EMPCNs, the primary limitation is that in geometric graphs, initial connectivity may not be available, and only a point cloud may be provided. While pre-domain knowledge, as discussed in Section 5.3, can be used to identify important higher-order structures, a promising research direction would be to develop methods to autonomously identify these structures based on the downstream task. Finally, although our method is proposed for cellular complexes, the same technique could be applied to higher-order path message passing, potentially resulting in a more expressive framework by incorporating different types of structures that are known to be useful for a particular downstream task.

### v3

@@ 5.1 N-Body System
As introduced by Kipf et al. (2018), the N-body system experiment involves tracking the trajectories of five charged particles in three-dimensional space over time. The objective is to predict the positions of all particles after 1,000 time steps, starting from their initial positions and velocities. To ensure a consistent comparison, we adopted the experimental setup as described by Satorras et al. (2021). While the system operates according to physics, with particles being attracted or repelled based on their charges, the task remains E(n) equivariant, maintaining these symmetries throughout the entire trajectory. This means that the predictions are consistent under Euclidean transformations, such as rotations and translations, ensuring the model respects these inherent physical symmetries.
@@
Task $\alpha$ $\Delta\varepsilon$ $\varepsilon_{\text{HOMO}}$ $\varepsilon_{\text{LUMO}}$ $\mu$ $C_{v}$ $G$ $H$ $R^{2}$ $U$ $U_{0}$ ZPVE Units (bohr3) (meV) (meV) (meV) (D) (cal/mol K) (meV) (meV) (bohr3) (meV) (meV) (meV) NMP .092 69 43 38 .030 .040 19 17 .180 20 20 1.50 SchNet .235 63 41 34 .033 .033 14 14 .073 19 14 1.70 Cormorant .085 61 34 38 .038 .026 20 21 .961 21 22 2.02 TFN .223 58 40 38 .064 .101 - - - - - - SE(3)-Tr. .142 53 35 33 .051 .054 - - - - - - DimeNet++ .043 32 24 19 .029 .023 7 6 .331 6 6 1.21 SphereNet .046 32 23 18 .026 .021 8 6 .292 7 6 1.21 PaiNN .045 45 27 20 .012 .024 7 6 .066 5 5 1.12 SEGNN .060 42 24 21 .023 .031 15 16 .660 12 15 1.62 MPSN .266 153 89 77 .101 .122 31 32 .887 33 33 3.02 EMPSN .066 37 25 20 .023 .024 6 9 .101 7 10 1.37 EGNN .071 48 29 25 .029 .031 12 12 .106 12 12 1.55 Decoupled EMPCN .063 40 27 22 .026 .027 10 10 .104 10 11 1.51 Gain 11% 17% 7% 12% 10% 13% 17% 17% 2% 17% 9% 3%
@@ 5.2 QM9
As indicated in Table 2, incorporating this additional ring information yielded an average improvement of approximately 10% over the baseline EGNN model. While one could potentially generate additional rings via the RadiusGraph or Vietoris-Rips complex, our experiment demonstrates that even with the initial graph’s topology, we can achieve significant performance gain without extra computational complexity. The reason for not applying the traditional EMPCNs in this experiment is because, in the fully connected setting, the number of rings becomes extremely large, and calculating the respective invariants is computationally infeasible.
@@ 5.4 Ablation Study
These findings highlight the robustness and efficiency of decoupled EMPCNs, illustrating that they are a more reliable choice for scenarios where data and computational resources are limited as is typical in real-life settings. Robustness and efficiency here refer to maintaining higher performance than EGNN across different data regimes and model complexities.
@@ 6 Conclusion
In summary, decoupled EMPCNs provide a scalable and expressive framework for higher-order message passing in geometric and topological graphs. Although this method is proposed for cellular complexes, the same techniques could be adapted to other higher-order path message passing methods, potentially enhancing the expressiveness by incorporating various structures beneficial for specific tasks. We believe that exploring the intersection between geometric and topological MPNNs is a crucial aspect of developing efficient and strongly generalizing models, especially in low-data settings.
@@
The main limitation of EMPCNs is the computational cost associated with the calculation of the introduced geometric invariants. Although it is possible to retain only those invariants relevant to a specific downstream task, an interesting direction for future research would be to pre-aggregate cell representations and compute the geometric invariants solely for key structures essential to the task. However, it is worth mentioning that these invariants are significantly cheaper to compute than the ones in methods using e.g. Clifford algebras.
For the decoupled EMPCNs, the primary limitation is that in geometric graphs, initial connectivity may not be available, and only a point cloud may be provided. While pre-domain knowledge, as discussed in Section 5.3, can be used to identify important higher-order structures, a promising research direction would be to develop methods to autonomously identify these structures based on the downstream task. In addition, as we just showed with the decoupled EMPCNs a generic template that can combine geometric and topological approaches, our goal was to beat the performance of the main architecture being used, which was EGNN in our case. Further experiments with more advanced architectures like SEGNN (Brandstetter et al., 2021) should be tested. Finally, although our method is proposed for cellular complexes, the same technique could be applied to higher-order path message passing, potentially resulting in a more expressive framework by incorporating different types of structures that are known to be useful for a particular downstream task.